OpenStudy (anonymous):

A 160 foot long piece of yarn is cut into 3 pieces. The second piece of yarn is four times as long as the first piece of yarn. The third piece of yarn is three times as long as the second piece of yarn. What is the length of the second piece to the nearest foot? Work is optional. I'm just stuck.

OpenStudy (anonymous):

If we say Length1= x, then we know Length2 = 4x and, Length3 = 3 x Length2 = 3 x 4x 12x Knowing this, we can say that the yarn (160ft) = Length1 + Length2 + Length3 So, 160 = x + 4x + 12x = 17x So, x =160/17 And Length2 = 4x = 4(160/17) = 640/17 = 37.647ft

OpenStudy (anonymous):

Hope that helps! :)

OpenStudy (anonymous):

YES!!! Thank you so much!!

OpenStudy (anonymous):

No problem! :)